Labeau Products, Ltd., of Perth, Australia, has $18,000 to invest. The company is trying to decide between two altemative uses for the funds as follows: Invest in Invest in Pro $18,000 18,000 $ 7,000 Investment required Annual cash inflows Single cash inflow at the end of 6 years Life of the project $ 41,000 6 years 6 years The companys discount rate is 17%. Click here to view Exhibit 11B-1 and Exhibit 11B-2, to determine the appropriate discount factor(s) using tables Required: a. Determine the net present values Project X Project Y Net Present Value b. Which alternative would you recommend that the company accept? O Project X O Project Y
Answer #1

a)

Project X , NPV = Annual inflow * PVIFA(dis rate , n) - Initial investment = 7000 * PVIFA(17,6)- 18000 = 7000*3.5892-18000 = $7124.40

Project Y , NPV = 6th Annum inflow * PVIF(dis rate , n) - Initial investment = 41000 * PVIF(17,6)- 18000 = 41000*0.3898 - 18000 = -$2018.20

b) Company should accept Project X.